Find the indefinite integral of the following functions
Y= 3t^2+2e^3t+1/t+ cos 3t
Y=x/2+3x^2
1
Expert's answer
2020-04-27T18:51:18-0400

"\u222b(3t^2+2e^{3t} +1\/t+cos3t)dt=3\u222bt^2 dt+2\u222be^{3t}dt +\u222bdt\/t+\u222bcos(3t)dt=" "=3\u22c5(1\/3) t^3+2\u22c5(1\/3) \u2147^{3t}+ln\u2061|t|+(1\/3) sin\u2061(3t)+C=t^3+(2\/3) \u2147^{3t}+ln\u2061|t|+(1\/3) sin(\u20613t)+C\\\\"


"\\2) ∫( x/2+3x^2)dx=(1/2) ∫xdx+3∫x^2 dx="

"=(1\/2)\u22c5(1\/2) x^2+3\u22c5(1\/3) x^3+C=(1\/4) x^2+x^3+C"